OpenStudy (anonymous):

find the value of the constant (c) that makes the function continuous. f(x)={x^2-c for x < 5 {4x+2c for x ≥ 5

OpenStudy (anonymous):

I think something like this should work, solve for c: \[\lim_{x \rightarrow 5} x^2 - c = 4(5) + 2c\]

OpenStudy (anonymous):

I figured it out actually. \[x ^{2}-c=4x+2c\] when x =5, so the answer is \[c=\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

but yeah thats the same thing as what you gave pretty much.